# Efficient Learning for Entropy-regularized Markov Decision Processes via Multilevel Monte Carlo

### v1

@@ 5.5 Analysis of the Bias
$\delta^{\mathrm{bias}}_{n,M,\mathbf{T}}\leq\gamma L\left\|Q^{0}_{n-1,M,\mathbf{T}}(S^{0}_{s,a},A^{0})-Q^{\star}(S^{0}_{s,a},A^{0})\right\|_{L^{2}}+\gamma\delta_{\mathbf{T}}(s,a),$
@@
$\delta^{\text{bias}}_{n,M,\mathbf{T}}$ $=\gamma\left|\mathbb{E}\left[T^{0}Q_{n-1,M,\mathbf{T}}^{0}(S^{0}_{s,a})\right]-\mathbb{E}\left[TQ^{\star}(S_{s,a}^{0})\right]\right|$
$\leq\gamma\left|\mathbb{E}\left[T^{0}Q_{n-1,M,\mathbf{T}}^{0}(S^{0}_{s,a})\right]-\mathbb{E}\left[T^{0}Q^{\star}(S^{0}_{s,a})\right]\right|+\gamma\left|\mathbb{E}\left[T^{0}Q^{\star}(S^{0}_{s,a})\right]-\mathbb{E}\left[TQ^{\star}(S^{0}_{s,a})\right]\right|$
$\leq\gamma L\mathbb{E}\left|Q^{0}_{n-1,M,\mathbf{T}}(S^{0}_{s,a},A^{(0,1)})-Q^{\star}(S^{0}_{s,a},A^{(0,1)})\right|+\gamma\delta_{\mathbf{T}}(s,a)$
$\leq\gamma L\left\|Q^{0}_{n-1,M,\mathbf{T}}(S^{0}_{s,a},A^{(0,1)})-Q^{\star}(S^{0}_{s,a},A^{(0,1)})\right\|_{L^{2}}+\gamma\delta_{\mathbf{T}}(s,a).$
@@ 5.6 Putting Everything Together: Global Error
$\leq\frac{(1+\rho_{n,M})\gamma\sigma_{\mathbf{T}}(s,a)}{\sqrt{M^{n}}}+\sum_{l=1}^{n-1}\frac{(1+\rho_{n,M})\gamma L}{\sqrt{M^{n-l}}}\left\|Q_{l,M,\mathbf{T}}^{(0,l)}(S_{s,a}^{0},A^{0})-Q_{l-1,M,\mathbf{T}}^{(0,-l)}(S_{s,a}^{0},A^{0})\right\|_{L^{2}}$
$+\gamma L\left\|Q^{0}_{n-1,M,\mathbf{T}}(S^{0}_{s,a},A^{(0,1)})-Q^{\star}(S^{0}_{s,a},A^{(0,1)})\right\|_{L^{2}}+\gamma\delta_{\mathbf{T}}(s,a),$
@@
$b_{1}$ $=\max\left(\|Q_{0}-Q^{\star}\|_{\infty},\tilde{\gamma}\|\sigma_{\mathbf{T}}\|_{\infty}\right)\geq 0,$
$b_{2}$ $=\gamma\|\delta_{\mathbf{T}}\|_{\infty}\geq 0,\quad b_{3}=\sqrt{M}\geq 0,$
$\lambda_{1}$ $=L(\tilde{\gamma}+\gamma\sqrt{M})\geq 0,$
$\lambda_{2}$ $=\tilde{\gamma}L(1+\sqrt{M})-\lambda_{1}=(\tilde{\gamma}-\gamma)\sqrt{M}\geq 0.$
@@
$\Lambda=\frac{1+L(\tilde{\gamma}+\gamma\sqrt{M})+\sqrt{\left(1+L(\tilde{\gamma}+\gamma\sqrt{M})\right)^{2}+4(\tilde{\gamma}-\gamma)\sqrt{M}}}{2},$
@@
$\frac{\Lambda}{\sqrt{M}}=\frac{\gamma L+\frac{1+\tilde{\gamma}L}{\sqrt{M}}+\sqrt{\left(\gamma L+\frac{1+\tilde{\gamma}L}{\sqrt{M}}\right)^{2}+4\frac{\tilde{\gamma}-\gamma}{\sqrt{M}}}}{2}\leq\gamma L+\frac{1+\tilde{\gamma}L}{\sqrt{M}}+\frac{\sqrt{\tilde{\gamma}-\gamma}}{M^{1/4}}<1,$ (5.4)
@@
$E_{n,M,\mathbf{T}}\leq\frac{3}{2}\Bigg{(}\max\left(\|Q_{0}-Q^{\star}\|_{\infty},\tilde{\gamma}\|\sigma_{\mathbf{T}}\|_{\infty}\right)\left[\gamma L+\frac{1+\tilde{\gamma}L}{\sqrt{M}}+\frac{\sqrt{\tilde{\gamma}-\gamma}}{M^{1/4}}\right]^{n}\\ +\gamma\|\delta_{\mathbf{T}}\|_{\infty}\frac{\sqrt{M}}{\sqrt{M}-\Lambda}\Bigg{)},$ (5.6)
@@ 5.7 Specializing with a Plain Monte Carlo Estimator
$E_{n,M,\mathbf{T}_{K}}\leq\frac{3}{2}\left(\max\left(\|Q_{0}-Q^{\star}\|_{\infty},2\gamma\|\sigma_{\mathbf{T}_{K}}\|_{\infty}\right)\left(\Lambda_{M}\right)^{n}+\frac{\gamma L^{\prime}}{\sqrt{K}}\frac{1}{1-\Lambda_{M}}\right),$
where $L^{\prime}\coloneqq\tau\left[L-1\right]$, and $\sigma_{\mathbf{T}_{K}}$ is defined in Theorem 2.2.
@@ 6.1 Complexity of the MLMCb Estimator
Observe that when using the plain Monte Carlo approximations $\mathbf{T}_{K}=(\hat{T}^{\theta}_{K})_{\theta\in\Theta}$, computing a realization of $\hat{T}^{\theta}_{K}Q(s)$ exactly requires $K$ samples from $\mu$ given by $A^{(\theta,1)},\\ A^{(\theta,2)},\ldots,A^{(\theta,K)}$. Let $\mathfrak{C}_{n,M,K}$ be the total number of independent random variables needed to compute a sample of $Q^{0}_{n,M,\mathbf{T}_{K}}(s,a)$. We have for $M\geq 1$
@@
$n_{\varepsilon}$ $=\left\lceil\frac{\log\varepsilon-\log D}{\log\Lambda_{M_{0}}}\right\rceil,\quad K_{\varepsilon}=\left\lceil 9\varepsilon^{-2}\left(\gamma L^{\prime}\frac{1}{1-\Lambda_{M_{0}}}\right)^{2}\right\rceil$
where $\Lambda_{M_{0}}=\gamma L+M_{0}^{-1/2}(1+2\gamma L)+M_{0}^{-1/4}\sqrt{\gamma}<1$. As such, $M_{0}$ is just a function of $\gamma,\tau,c_{\min},c_{\max}$. Let $\tilde{C}(M_{0})\coloneqq(1-\Lambda_{M_{0}})^{-1}9\gamma L^{\prime}$. With that choice of $M,n,K$, the complexity bound (6.3) can be written as
$\mathfrak{C}_{n_{0},M_{0},K_{0}}\leq 2^{-l/l^{\prime}+3}(\tilde{C}(M_{0})+1)^{-l/l^{\prime}+2}M_{0}^{-l/l^{\prime}+1}\varepsilon^{-2\log(\varepsilon)/l^{\prime}+\frac{2l+\log 2+\log M_{0}+\log(\tilde{C}(M_{0})+1)}{l^{\prime}}-4},$
@@
$n_{0}\leq\frac{\log\varepsilon-\log D}{\log\Lambda_{M_{0}}}+1,\quad K_{0}\leq\varepsilon^{-2}\left(\tilde{C}(M_{0})+1\right),$
@@ 6.2 Complexity of the MLMCu Estimator
The sample complexity now becomes a random variable due to the random sample size of $\tilde{T}^{\theta}$. Recall that $\tilde{K}+1$ is a geometric random variable with parameter $r\in(1/2,3/4)$ such that $2^{\tilde{K}+1}$ corresponds to the number of i.i.d. copies of $\mu$ needed to compute a realization of $\tilde{T}^{\theta}Q(s)$ for any fixed $\theta,Q,s$. In particular, we need to handle the recursion in the cost carefully. The analysis performed in (6.1) becomes
$\mathfrak{C}^{\theta}_{n,M}=\sum_{i=1}^{M^{n}}(2^{\tilde{K}^{(\theta,0,i)}+2})+\sum_{l=1}^{n-1}\sum_{i=1}^{M^{n-l}}(1+2^{\tilde{K}^{(\theta,l,i)}+1}(\mathfrak{C}^{(\theta,l,i)}_{l,M}+\mathfrak{C}^{(\theta,-l,i)}_{l-1,M}+1)).$
Finally, the random variable $K^{(\theta,l,i)}$ is independent of both $\mathfrak{C}^{(\theta,l,i)}_{l,M}$ and $\mathfrak{C}^{(\theta,-l,i)}_{l-1,M}$. In particular, letting $\mathfrak{C}_{n,M}=\mathbb{E}\mathfrak{C}^{\theta}_{n,M}$ and $K=\mathbb{E}2^{\tilde{K}^{\theta}+1}$, we recover (6.1). Now, $K$ does not depend on $n$, hence we get a polynomial bound in expectation. Notice how the choice of the parameter $r>1/2$ is important here, since it ensures that $K=\mathbb{E}2^{\tilde{K}+1}=\sum_{n\geq 0}2^{n+1}p(n)<\infty$.
@@
$\mathfrak{C}_{n,M}^{\theta}=\sum_{i=1}^{M^{n}}\left(1+2^{\tilde{K}^{(\theta,0,i)}+1}+1\right)+\sum_{l=1}^{n-1}\sum_{i=1}^{M^{n-l}}\left(1+(2^{\tilde{K}^{(\theta,l,i)}+1}+1)(\mathfrak{C}^{(\theta,l,i)}_{l,M}+\mathfrak{C}^{(\theta,-l,i)}_{l-1,M}+1)\right).$
@@
$+\sum_{l=1}^{n-1}\sum_{i=1}^{M^{n-l}}\left(1+\mathbb{E}\left[2^{\tilde{K}^{(\theta,l,i)}+1}+1\right](\mathbb{E}\mathfrak{C}^{(\theta,l,i)}_{l,M}+\mathbb{E}\mathfrak{C}^{(\theta,-l,i)}_{l-1,M}+1)\right).$
@@
We can apply Gronwall’s inequality from Lemma A.1 to get (6.2), i.e
@@ 7 Conclusion and Future Work
We begin by proving that a simple iterative MC algorithm achieves quasi-polynomial sample complexity. To improve performance, we introduce a general framework for constructing multilevel Monte Carlo (MLMC) estimators which combine fixed-point iteration, MLMC techniques, and a suitable stochastic approximation of the Bellman operator. We quantify the $L^{2}$ error of the MLMC estimator in terms of the properties of the chosen approximate Bellman operator. Building on this error analysis, we show that using a biased plain MC estimate for the Bellman operator results in an MLMC estimator that achieves a quadratic reduction in sample complexity compared to the simple iterative MC estimator, though it still suffers from quasi-polynomial complexity due to the inherent bias. We then adapt debiasing techniques from [4] to construct an unbiased randomized multilevel approximation of the Bellman operator. The resulting MLMC estimator achieves polynomial sample complexity in expectation, providing the first polynomial-time estimator for general action spaces. Along the way, we also prove the Lipschitz continuity of the Blanchet–Glynn estimator with respect to the input random variable, which is a result of independent interest.
A natural extension of our work would be to analyze how to sample from the policy derived from the estimated Q-function. Additionally, the proposed estimators could be adapted to partially observable MDPs or mean-field MDPs. Finally, combining the MLMC with other reinforcement learning methods is a worthwhile direction for future research.
8 Acknowledgments
MM is supported by the EPSRC Centre for Doctoral Training in Mathematics of Random Systems: Analysis, Modelling and Simulation (EP/S023925/1). The authors are grateful for comments and suggestions made by Mike Giles and Justin Sirignano.
@@ Appendix B Proofs of Technical Lemmas
$S_{E}Q^{\prime}(2^{N})$ $=\frac{1}{2^{N}}\sum_{k=1}^{2^{N}}\exp\left(-Q^{\prime}(s,A_{2k})/\tau\right),$
$S_{O}Q^{\prime}(2^{N})$ $=\frac{1}{2^{N}}\sum_{k=1}^{2^{N}}\exp\left(-Q^{\prime}(s,A_{2k-1})/\tau\right),$
($E/O$ making a reference to the even / odd indices used in the sum). Notice that for any $Q^{\prime}$, $SQ^{\prime}(2^{N+1})=\frac{1}{2}(S_{E}Q^{\prime}(2^{N})+S_{O}Q^{\prime}(2^{N}))$. We now examine the difference $\Delta_{N}^{\theta}Q_{1}(s)-\Delta_{N}^{\theta}Q_{2}(s)$ which we decompose in 3 terms
@@
where $D_{N}=g(SQ_{1}(2^{N+1}))-g(SQ_{2}(2^{N+1}))$, $D_{E}(N)=g(S_{E}Q_{1}(2^{N}))-g(S_{E}Q_{2}(2^{N}))$, and $D_{O}(N)$ is defined likewise. Since the sums at which $g$ is evaluated are lower-bounded by $A=e^{-\beta/\tau}$ and upper-bounded by $B=e^{-\alpha/\tau}$, we can apply Lemma 5.7 to
$a_{N}=S_{E}Q_{1}(2^{N}),\quad b_{N}=S_{O}Q_{1}(2^{N}),\quad a^{\prime}_{N}=S_{E}Q_{2}(2^{N}),\quad b^{\prime}_{N}=S_{O}Q_{2}(2^{N}),$
which gives an upper bound like (5.8) with the constant given by $\tau C_{2}(e^{-\beta/\tau},e^{-\alpha/\tau})$. Notice that $m=\int_{\mathcal{A}}e^{-Q_{1}(s,a)/\tau}\mu({\mathrm{d}}a)$, and $m^{\prime}=\int_{\mathcal{A}}e^{-Q_{2}(s,a)/\tau}\mu({\mathrm{d}}a)$. We now show that each of the terms in the upper bound (5.8) can be bounded in squared $L^{2}$ norm by
$\frac{D_{(\alpha,\tau)}}{\tau^{2}2^{2N}}\int_{\mathcal{A}}|Q_{1}(s,a)-Q_{2}(s,a)|^{2}\mu({\mathrm{d}}a),$
@@
where the sum is taken over all $k_{1},k_{2},k_{3},k_{4},k_{5},k_{6}\in\{1,\cdots,2^{N}\}$. Let $K=2^{N}$, we now aim to prove that at most $K^{4}$ (up to a numerical factor) of the terms in the sum are non zero. Notice that
$\mathbb{E}|q_{k_{1}}-r_{k_{1}}||q_{k_{2}}-r_{k_{2}}|\prod_{i=3}^{6}(e^{-q_{k_{i}}/\tau}-m)\neq 0\quad\text{ implies }\quad\forall i\in\{3,4,5,6\},\exists j\neq i,k_{i}=k_{j}.$
It suffices to find an upper-bound of the cardinality of the set
@@
$\leq\frac{D_{(\alpha,\tau)}}{\tau^{2}2^{2N}}\mathbb{E}|q_{1}-r_{1}|^{2}$
$=\frac{D_{(\alpha,\tau)}}{\tau^{2}2^{2N}}\int_{\mathcal{A}}|Q_{1}(s,a)-Q_{2}(s,a)|^{2}\mu({\mathrm{d}}a),$

### v3

@@ 5.5 Analysis of the Bias
$\delta^{\mathrm{bias}}_{n,M,\mathbf{T}}\leq\gamma L\left\|Q^{0}_{n-1,M,\mathbf{T}}\left(S^{0}_{s,a},A^{0}\right)-Q^{\star}\left(S^{0}_{s,a},A^{0}\right)\right\|_{L^{2}}+\gamma\delta_{\mathbf{T}}(s,a),$
@@
$\delta^{\text{bias}}_{n,M,\mathbf{T}}$ $=\gamma\left|\mathbb{E}\left[T^{0}Q_{n-1,M,\mathbf{T}}^{0}\left(S^{0}_{s,a}\right)\right]-\mathbb{E}\left[TQ^{\star}\left(S_{s,a}^{0}\right)\right]\right|$
$\leq\gamma\left|\mathbb{E}\left[T^{0}Q_{n-1,M,\mathbf{T}}^{0}\left(S^{0}_{s,a}\right)\right]-\mathbb{E}\left[T^{0}Q^{\star}\left(S^{0}_{s,a}\right)\right]\right|$
$+\gamma\left|\mathbb{E}\left[T^{0}Q^{\star}\left(S^{0}_{s,a}\right)\right]-\mathbb{E}\left[TQ^{\star}\left(S^{0}_{s,a}\right)\right]\right|$
$\leq\gamma L\mathbb{E}\left|Q^{0}_{n-1,M,\mathbf{T}}\left(S^{0}_{s,a},A^{(0,1)}\right)-Q^{\star}\left(S^{0}_{s,a},A^{(0,1)}\right)\right|+\gamma\delta_{\mathbf{T}}(s,a)$
$\leq\gamma L\left\|Q^{0}_{n-1,M,\mathbf{T}}\left(S^{0}_{s,a},A^{(0,1)}\right)-Q^{\star}\left(S^{0}_{s,a},A^{(0,1)}\right)\right\|_{L^{2}}+\gamma\delta_{\mathbf{T}}(s,a).$
@@ 5.6 Putting Everything Together: Global Error
$\leq\frac{(1+\rho_{n,M})\gamma\sigma_{\mathbf{T}}(s,a)}{\sqrt{M^{n}}}+\sum_{l=1}^{n-1}\frac{(1+\rho_{n,M})\gamma L}{\sqrt{M^{n-l}}}\left\|Q_{l,M,\mathbf{T}}^{(0,l)}\left(S_{s,a}^{0},A^{0}\right)-Q_{l-1,M,\mathbf{T}}^{(0,-l)}\left(S_{s,a}^{0},A^{0}\right)\right\|_{L^{2}}$
$+\gamma L\left\|Q^{0}_{n-1,M,\mathbf{T}}\left(S^{0}_{s,a},A^{(0,1)}\right)-Q^{\star}\left(S^{0}_{s,a},A^{(0,1)}\right)\right\|_{L^{2}}+\gamma\delta_{\mathbf{T}}(s,a),$
@@
$\quad b_{1}$ $=\max\left(\|Q_{0}-Q^{\star}\|_{\infty},\tilde{\gamma}\|\sigma_{\mathbf{T}}\|_{\infty}\right)\geq 0,\quad b_{2}=\gamma\|\delta_{\mathbf{T}}\|_{\infty}\geq 0,\quad b_{3}=\sqrt{M}\geq 0,$
$\lambda_{1}$ $=L(\tilde{\gamma}+\gamma\sqrt{M})\geq 0,\quad\lambda_{2}=\tilde{\gamma}L(1+\sqrt{M})-\lambda_{1}=(\tilde{\gamma}-\gamma)\sqrt{M}\geq 0.$
@@
$\Lambda=\frac{1}{2}\left(1+L(\tilde{\gamma}+\gamma\sqrt{M})+\sqrt{\left(1+L(\tilde{\gamma}+\gamma\sqrt{M})\right)^{2}+4(\tilde{\gamma}-\gamma)\sqrt{M}}\right),$
@@
$\frac{\Lambda}{\sqrt{M}}=\frac{1}{2}\left(\gamma L+\frac{1+\tilde{\gamma}L}{\sqrt{M}}+\sqrt{\left(\gamma L+\frac{1+\tilde{\gamma}L}{\sqrt{M}}\right)^{2}+4\frac{\tilde{\gamma}-\gamma}{\sqrt{M}}}\right)\leq\gamma L+\frac{1+\tilde{\gamma}L}{\sqrt{M}}+\frac{\sqrt{\tilde{\gamma}-\gamma}}{M^{1/4}}<1,$ (5.4)
@@
$E_{n,M,\mathbf{T}}\leq\\ \frac{3}{2}\left(\max\left(\|Q_{0}-Q^{\star}\|_{\infty},\tilde{\gamma}\|\sigma_{\mathbf{T}}\|_{\infty}\right)\left[\gamma L+\frac{1+\tilde{\gamma}L}{\sqrt{M}}+\frac{\sqrt{\tilde{\gamma}-\gamma}}{M^{1/4}}\right]^{n}+\gamma\|\delta_{\mathbf{T}}\|_{\infty}\frac{\sqrt{M}}{\sqrt{M}-\Lambda}\right),$ (5.6)
@@ 5.7 Specializing with a Plain Monte Carlo Estimator
$E_{n,M,\mathbf{T}_{K}}\leq\frac{3}{2}\left(\max\left(\|Q_{0}-Q^{\star}\|_{\infty},2\gamma\|\sigma_{\mathbf{T}_{K}}\|_{\infty}\right)\left(\Lambda_{M}\right)^{n}+\frac{\gamma(L^{\prime})^{2}}{2\tau K}\frac{1}{1-\Lambda_{M}}\right),$
where $L^{\prime}\coloneqq\tau(L-1)$, and $\sigma_{\mathbf{T}_{K}}$ is defined in Theorem 2.2.
@@ 6.1 Complexity of the MLMCb Estimator
Observe that when using the plain Monte Carlo approximations $\mathbf{T}_{K}=(\hat{T}^{\theta}_{K})_{\theta\in\Theta}$, computing a realization of $\hat{T}^{\theta}_{K}Q(s)$ exactly requires $K$ samples from $\mu$ given by $A^{(\theta,1)},\allowbreak A^{(\theta,2)},\ldots,A^{(\theta,K)}$. Let $\mathfrak{C}_{n,M,K}$ be the total number of independent random variables needed to compute a sample of $Q^{0}_{n,M,\mathbf{T}_{K}}(s,a)$. We have for $M\geq 1$
@@
$n_{\varepsilon}$ $=\left\lceil\frac{\log\varepsilon-\log D}{\log\Lambda_{M_{0}}}\right\rceil,\quad K_{\varepsilon}=\left\lceil 3\frac{\gamma(L^{\prime})^{2}}{2\tau(1-\Lambda_{M_{0}})\varepsilon}\right\rceil$
where $\Lambda_{M_{0}}=\gamma L+M_{0}^{-1/2}(1+2\gamma L)+M_{0}^{-1/4}\sqrt{\gamma}<1$. As such, $M_{0}$ is just a function of $\gamma,\tau,c_{\min},c_{\max}$. Let $\tilde{C}(M_{0})\coloneqq(2\tau(1-\Lambda_{M_{0}}))^{-1}3\gamma(L^{\prime})^{2}$. With that choice of $M,n,K$, the complexity bound (6.3) can be written as
$\mathfrak{C}_{n_{0},M_{0},K_{0}}\leq 2^{-l/l^{\prime}+3}(\tilde{C}(M_{0})+1)^{-l/l^{\prime}+2}M_{0}^{-l/l^{\prime}+1}\varepsilon^{-\log(\varepsilon)/l^{\prime}+\frac{2l+\log 2+\log M_{0}+\log(\tilde{C}(M_{0})+1)}{l^{\prime}}-4},$
@@
$n_{\varepsilon}\leq\frac{\log\varepsilon-\log D}{\log\Lambda_{M_{0}}}+1,\quad K_{\varepsilon}\leq\varepsilon^{-1}\left(\tilde{C}(M_{0})+1\right),$
@@ 6.2 Complexity of the MLMCu Estimator
The sample complexity now becomes a random variable due to the random sample size of $\tilde{T}^{\theta}$. Recall that $\tilde{K}+1$ is a geometric random variable with parameter $r\in(1/2,3/4)$ such that $2^{\tilde{K}+1}$ corresponds to the number of i.i.d. copies of $\mu$ needed to compute a realization of $\tilde{T}^{\theta}Q(s)$ for any fixed $\theta,Q,s$. In particular, we need to handle the recursion in the cost carefully, as the sample complexity $\mathfrak{C}^{\theta}_{n,M}$ of the MLMCu estimator $Q^{\theta}_{n,M}$ is now stochastic. Notice that the random variable $K^{(\theta,l,i)}$ is independent of both $\mathfrak{C}^{(\theta,l,i)}_{l,M}$ and $\mathfrak{C}^{(\theta,-l,i)}_{l-1,M}$. In particular, letting $\mathfrak{C}_{n,M}=\mathbb{E}\mathfrak{C}^{\theta}_{n,M}$ and $K=\mathbb{E}2^{\tilde{K}^{\theta}+1}+1$, we recover (6.1). Now, $K$ does not depend on $n$, hence we get a polynomial bound in expectation. Notice how the choice of the parameter $r>1/2$ is important here, since it ensures that $K=\mathbb{E}2^{\tilde{K}+1}=\sum_{n\geq 0}2^{n+1}p(n)<\infty$.
@@
$\mathfrak{C}_{n,M}^{\theta}=\sum_{i=1}^{M^{n}}\left(1+2^{\tilde{K}^{(\theta,0,i)}+1}+1\right)+\sum_{l=1}^{n-1}\sum_{i=1}^{M^{n-l}}\left(1+\left(2^{\tilde{K}^{(\theta,l,i)}+1}+1\right)\left(\mathfrak{C}^{(\theta,l,i)}_{l,M}+\mathfrak{C}^{(\theta,-l,i)}_{l-1,M}+1\right)\right).$
@@
$+\sum_{l=1}^{n-1}\sum_{i=1}^{M^{n-l}}\left(1+\mathbb{E}\left[2^{\tilde{K}^{(\theta,l,i)}+1}+1\right]\left(\mathbb{E}\mathfrak{C}^{(\theta,l,i)}_{l,M}+\mathbb{E}\mathfrak{C}^{(\theta,-l,i)}_{l-1,M}+1\right)\right).$
@@
We can apply Gronwall’s inequality from Lemma A.1 to get (6.2), i.e.,
@@ 7 Conclusion and Future Work
We begin by proving that a simple iterative MC algorithm achieves quasi-polynomial sample complexity. To improve performance, we introduce a general framework for constructing multilevel Monte Carlo (MLMC) estimators which combine fixed-point iteration, MLMC techniques, and a suitable stochastic approximation of the Bellman operator. We quantify the $L^{2}$ error of the MLMC estimator in terms of the properties of the chosen approximate Bellman operator. Building on this error analysis, we show that using a biased plain MC estimate for the Bellman operator results in an MLMC estimator that achieves a cubic reduction in sample complexity compared to the simple iterative MC estimator, though it still suffers from quasi-polynomial complexity due to the inherent bias. We then adapt a debiasing technique from [4] to construct an unbiased randomized multilevel approximation of the Bellman operator. The resulting MLMC estimator achieves polynomial sample complexity in expectation, providing the first polynomial-time estimator for general action spaces. Along the way, we also prove the Lipschitz continuity of the Blanchet–Glynn estimator with respect to the input random variable, which is a result of independent interest.
A natural extension of this work is to investigate efficient sampling strategies from the policy induced by the estimated Q-function. Moreover, the proposed estimators could be generalized to more complex settings, such as partially observable or mean-field MDPs. Finally, integrating MLMC techniques with other (deep) reinforcement learning approaches presents a promising direction for future research.
8 Acknowledgements and Disclosure of Funding
MM is supported by the EPSRC Centre for Doctoral Training in Mathematics of Random Systems: Analysis, Modelling and Simulation (EP/S023925/1). CR is partially supported by EPSRC grant EP/Y028872/1. The authors are grateful for comments and suggestions made by Mike Giles and Justin Sirignano.
@@ Appendix B Proofs of Technical Lemmas
$S_{E}Q^{\prime}(2^{N})$ $=\frac{1}{2^{N}}\sum_{k=1}^{2^{N}}\exp\left(-Q^{\prime}(s,A_{2k})/\tau\right),\quad S_{O}Q^{\prime}(2^{N})=\frac{1}{2^{N}}\sum_{k=1}^{2^{N}}\exp\left(-Q^{\prime}(s,A_{2k-1})/\tau\right),$
$E/O$ referencing the even / odd indices used in the sum. Notice that for any $Q^{\prime}$, $SQ^{\prime}(2^{N+1})=\frac{1}{2}(S_{E}Q^{\prime}(2^{N})+S_{O}Q^{\prime}(2^{N}))$. We now examine the difference $\Delta_{N}^{\theta}Q_{1}(s)-\Delta_{N}^{\theta}Q_{2}(s)$ which we decompose in 3 terms
@@
where $D_{N}=g(SQ_{1}(2^{N+1}))-g(SQ_{2}(2^{N+1}))$, $D_{E}(N)=g(S_{E}Q_{1}(2^{N}))-g(S_{E}Q_{2}(2^{N}))$, and $D_{O}(N)$ is defined likewise. Since the sums at which $g$ is evaluated are lower-bounded by $A=e^{-\beta/\tau}$ and upper-bounded by $B=e^{-\alpha/\tau}$, we can apply Lemma 5.7 to $a_{N}=S_{E}Q_{1}(2^{N}),\quad b_{N}=S_{O}Q_{1}(2^{N}),\quad a^{\prime}_{N}=S_{E}Q_{2}(2^{N}),\quad b^{\prime}_{N}=S_{O}Q_{2}(2^{N})$, which gives an upper bound like (5.8) with the constant given by $\tau C_{2}(e^{-\beta/\tau},e^{-\alpha/\tau})$. Notice that $m=\int_{\mathcal{A}}e^{-Q_{1}(s,a)/\tau}\mu({\mathrm{d}}a)$ and $\quad m^{\prime}=\int_{\mathcal{A}}e^{-Q_{2}(s,a)/\tau}\mu({\mathrm{d}}a)$. We now show that each of the terms in the upper bound (5.8) can be bounded in squared $L^{2}$ norm by
$\tau^{-2}2^{-2N}D_{(\alpha,\tau)}\int_{\mathcal{A}}|Q_{1}(s,a)-Q_{2}(s,a)|^{2}\mu({\mathrm{d}}a),$
@@
where the sum is taken over all $k_{1},k_{2},k_{3},k_{4},k_{5},k_{6}\in\{1,\cdots,2^{N}\}$. Let $K=2^{N}$, we now aim to prove that at most $K^{4}$ (up to a numerical factor) of the terms in the sum are non zero. Notice that the non-zero terms correspond to $6$-tuples $(k_{1},\ldots,k_{6})$ such that for all $i\in\{3,4,5,6\}$, there exists $j\neq i,k_{i}=k_{j}$. It suffices to find an upper-bound of the cardinality of the set
@@
$\leq\frac{D_{(\alpha,\tau)}}{\tau^{2}2^{2N}}\mathbb{E}|q_{1}-r_{1}|^{2}=\frac{D_{(\alpha,\tau)}}{\tau^{2}2^{2N}}\int_{\mathcal{A}}|Q_{1}(s,a)-Q_{2}(s,a)|^{2}\mu({\mathrm{d}}a),$